slope=0; (4,-2) how do i put this in slope int form?
Answer by benazir.sj@gmail.com(109)

slope point form is
(y-y1)=m(x-x1)
m =0, x1=4 and y1 =-2
[y-(-2)] =0[x-4]
y+2 =0
the slope intercept form is
y=mx+b
ie,
y+2 =0
y=-2